OpenStudy (anonymous):

last problem i dont get: 3/n+1 = 4/n+4 last problem i dont get: 3/n+1 = 4/n+4 @Mathematics

OpenStudy (anonymous):

Cross Multiply

OpenStudy (anonymous):

3n + 12 = 4n +4

OpenStudy (anonymous):

is after you cross multiply...now solve for n

OpenStudy (anonymous):

8 = n

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

nono, n=8

OpenStudy (anonymous):

4n - 3n = n 12-4 = 8

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

zzzz

OpenStudy (turingtest):

\[{3\over n+1} = {4\over n+4}\rightarrow{3(n+4)\over (n+1)(n+4)}={4(n+1)\over (n+1)(n+4)}\]\[3n+12=4n+4\rightarrow n=8\]

OpenStudy (anonymous):

Ill give you a medal because thats a lot of typin' lol

OpenStudy (anonymous):

And.. the senior class is taking a trip to an amusement park. They received a special deal where for every 3 tickets they purchased they received one free ticket. 3 tickets cost $53.25. The total purchase of tickets cost; $1384.50. How many tickets did they receive?